# Analytical coarse grained potential parameterization by Reinforcement Learning for anisotropic cellulose

### v2

@@ 2.3 Boltzmann inversion for harmonic potential
However, the $k$ derived must be reparameterized to represent the mechanical performance[59, 60, 49] of the coarse-grained models with fewer degrees of freedom. The effective $k$ is $k_{0}$ times of $\kappa$:
@@ 2.4 Reinforcement Learning and degenerate RL
In this study, a degenerate version of RL was leveraged in a one-shot diagram. There is only one step in a learning episode, and only one reward is sampled, because the performance of a set of coefficients is independent of that of the others. Therefore, the reward can be simplified as
@@
The RL agent used in this study was driven by the Soft Actor Critic (SAC) algorithm[62]. SAC is the state-of-the-art algorithm recommended for continuous action space by the doc of OpenAI Stable-Baselines3[63]. The key feature is the reconciliation between the expected reward and entropy, which implies that stochasticity is included in the SAC policy to a certain extent. SAC is a gradient-free and model-free method that does not rely on a specific problem environment, and can address strongly nonlinear problems after sufficient training. Detailed information on the SAC algorithm can be obtained in the doc of OpenAI Stable-Baselines3[63].
During training, the action consists of 17 normalized coefficients ranging from [-1,1] because most RL algorithms rely on a standard normal distribution[63]. actions will be transformed linearly into physical space corresponding to 3 BD $k$, 1 HB $\epsilon$, 1 HB $\sigma$, 6 LJ $\epsilon$ and 6 LJ $\sigma$. The reward function is a critical factor that guides the learning of the agent, and is defined by a series of matching degrees. The matching degree $m$ for a specific performance is defined as
@@
. where $y_{0}$ and $y_{1}$ are the reference and sampled values, respectively. The reward function is the sum of a series of matching degrees, except for the transverse strength $M_{s}$ and toughness $M_{t}$,
$reward=\sum M_{i}+w\cdot{\rm min}(M_{s},M_{t})$ (17)
@@ 2.5 Simulation setup
For both all atom (AA) and coarse-grained (CG) simulations, three models in different arrangement directions under the NPT ensemble at 0.1 MPa with a stretch load in the vertical direction were simulated. All reference AA simulations were performed using GROMACS[65, 66] and CHARMM36[67, 68] force field (force field files were generated using CHARMM-GUI tools[69, 70]). All the CG simulations were performed using LAMMPS[54] with a self-defined force field. Referring to the performance-size curves obtained by Sinko et al.[71], we used a CNC model comprising 144 cellulose chains. Fragile vertical and horizontal models with 72 cellulose chains were used to reduce the computational cost of the simulations during training. Vertical stretches were implemented using cell deformation loads at a constant speed 10.0 nm/ns. 100 replica AA simulations with different stretch speed ranging from 0.1 nm/ns to 40.0 nm/ns are performed. The AA simulation results confirmed that the sampled strength and toughness were stable within this speed range (with deviations of less than 15%), as shown in Figure SP9. Figure LABEL:fig:aa_transverse_stretch_performance_collection shows the stress-strain curves of the 100 replica simulations.
@@ 2.6 Training procedures and statistics-guided reduction
Before training, the equilibrium geometry parameters (bond length $l^{e}$/,angle size $\theta^{e}$/, and dihedral size $\chi^{e}$) were derived from the mapped AA trajectories. The estimated force constants for the bonded harmonic interactions were determined using the same trajectories and Boltzmann inversion method. Three rescale factors were required for the bonds, angles, and improper dihedrals in this model. Although rescaled, the relative ratios of the force constants were maintained (e.g., the force constants for improper dihedrals were rescaled with the same factor). As previously mentioned, there are 17 coefficients, where three are BD coefficients and 14 are NB coefficients. To simplify the problem and reduce computational cost, three training stages were designed (Figure 4). The first stage was BD training for three BD $k$ only, ignoring all NB interactions and targeting the BD properties (axial elastic modulus and polymer stiffness). We cannot determine the influence of NB interactions in the first stage, and we assume that they are not critical for the BD properties. The second stage was NB training for 14 NB coefficients with the BD coefficients obtained from the first stage, targeting the BD and NB properties (transverse strength and toughness in the three characteristic directions). The third stage was preserved only if the performance in the second stage was unsatisfactory; all properties were targeted to parameterize 17 coefficients at every turn.
@@
Figure 4: Training procedures. The BD properties (axial elastic modulus and polymer stiffness) and NB properties (transverse strength and toughness in characteristic directions) were the target attributes. Before training, the equilibrium bonded geometry parameters and force constants of the BD harmonic potentials were estimated from the mapped AA trajectories using the Boltzmann inversion method. In the first stage, the force constants were scaled by the RL agent in the BD training to reproduce the BD properties. In BD training, all NB interactions were ignored because we assumed that the NB interactions were not significant in this case. Given both BD and NB properties, NB training using the previously obtained BD coefficients was performed to match both BD and NB properties. However, only one batch of simulations was performed for the target properties during training. Thus, the transverse stretch performance data were manually verified through replica simulations to confirm the statistical agreement. Multiple sets of coefficients were collected to reduce the number of independent coefficients for the computational cost and physical explainability. The results confirm that our assumption of the BD properties is reasonable. Therefore, the coupled training was unnecessary.
@@ 3.1 Coefficients
Figure 5: Training convergence and statistics-guided reduction. (a) Reward and its moving average in the first and (b) last training loops. Each training loop contained 1024 training steps. The subsequent training loops with higher and more stable rewards were based on former training loops. (c) Average and maximum rewards for different training loops. The average and maximum rewards confirmed the convergence. After training, the best coefficients for the different training loops were collected to perform statistics-guided reduction. The first reduction was inferred from the geometric characteristics of the CG models, and a constraint of $\sigma_{p11}=\sigma_{p22}=\sigma_{p33}$ was applied (pmn presents a pair of CLm and CLn), which is also helpful for fragile behaviors in the vertical and horizontal directions. (d) The second reduction was in the distance coefficients using statistics. $\sigma$ except for $\sigma_{p23}$ oscillated around its equilibrium distance $d^{e}$ whereas an obvious increase of $\sigma_{p23}$ was observed. Then only $\sigma_{p11}$ and $\sigma_{p23}$ are kept as independent distance coefficients: $\sigma_{p12}=\frac{d^{e}_{p12}}{d^{e}_{p11}}\sigma_{p11},\sigma_{p13}=\frac{d^{e}_{p13}}{d^{e}_{p11}}\sigma_{p11}$. (e) The third reduction in energy coefficients was statistically determined. After the first and second reduction, we observe that $\epsilon_{p23}$ is usually the smallest energy coefficients, which may correspond to the $\sigma{p23}$ enlargement. Then only $\epsilon_{p23}$ and $\epsilon_{p23}$ are kept as independent energy coefficient: $\epsilon_{p11}=\epsilon_{p22}=\epsilon_{p33}=\epsilon_{p12}=\epsilon_{p13}$.
The coefficients were trained using several training loops, each containing 1024 training steps. When a training loop was completed, the information of the agent was saved and loaded for subsequent training. Referring to the first training loop, the last training loop achieved a higher reward much faster and more stably (Figure 5a). The average and maximum rewards of the different training loops were compared (Figure 5b) to confirm the convergence of training.
@@
where pmnpmn is the pair of the CLm and CLn. The structures of the CG models with lower degrees of freedom were more orderly. Thus, the equilibrium distances of the same types should be close in order to represent the laminar structure. In other words, the equilibrium distance of CL1-CL1, CL2-CL2 and CL3-CL3 ought to be nearly the same, whereas in the mapped AA trajectories, they are not (Figure SP9). The equilibriumequilibrium distancesdistances of the ideal crystal structure were selected as the estimated values of the distance coefficients ($\sigma$) for the pairs. These equilibriumequilibrium distancesdistances do not correspond to the well-relaxed structures or characteristic distances of the radial distribution function. The first physics-informed restraint was also beneficial for reproducing fragile fractures in vertical and horizontal models.
@@
bond 1 64.06 5.270 5.270 5.258
2 82.42 2.493 2.493 2.473
3 112.3 2.066 2.066 2.061
kk(kcal/mol) $\theta_{0}(^{\circ})$ $\theta^{e}_{\rm AA}(^{\circ}))$ $\theta^{e}_{\rm CG}(^{\circ}))$
angle 1 30.97 163.5 163.5 162.3
2 38.79 74.2 74.2 74.7
3 30.52 93.2 93.2 94.0
4 27.28 90.0 90.0 89.2
5 25.97 102.0 102.0 101.4
6 22.42 166.3 166.3 165.1
kk(kcal/mol) $\chi_{0}(^{\circ})$ $\chi^{e}_{\rm AA}(^{\circ}))$ $\chi^{e}_{\rm CG}(^{\circ}))$
improper 1 0.1702 178.0 178.0 175.6
2 0.1555 2.3 2.3 2.1
3 0.1698 2.9 2.9 2.1
@@
type pair $\epsilon$(kcal/mol/Å) $\sigma$(Å) $d^{e}_{\rm AA}$(Å) $d^{e}_{\rm CG}$(Å)
@@ 3.2 Bonded properties
For the BD interactions and validation of the BD coefficients, we considered the axial elastic modulus and polymer stiffness of a single cellulose chain, as listed in Table 3 and shown in Figure 6. Axial stretching was performed in an xyz-periodic cell with a deformation load along the chains, which is a basic mechanical property. The simulation of polymer stiffness is a single chain centered in a cell. For axial stretch, Poisson effect was presented in CG model and elastic modulus of AA and CG are 133.5 GPa and 130.3 GPa respectively, where reference modulus is 120-140 GPa[13, 14] from experiments and 110-200 GPa[16, 38, 17] from simulations. For polymer stiffness, persistence length of AA and CG range from 9.7-9.8 nm and 8.9-10.2 nm with reference data of 6-25 nm from experiments[72] and simulation[73]. Notably, the polymer stiffness data were not stable for the AA simulations because of the intertwining in some cases. Therefore, persistence lengths and end-to-end distances of different lengths (10, 20, and 30 residues) are listed. When considering polymer stiffness, the joint O atoms between two cellulose residues in the AA models and CL1 beads on the backbone in the CG models were node particles to be considered. The reference data for training were based on AA information only, and the mapped AA trajectories were compared with the CG trajectories after training. In this study, all effective $k$ values were smaller than their estimations based on distributions (Table SP2), particularly for angles and dihedrals, to reproduce the polymer stiffness performance in a system with fewer degrees of freedom (Table SP2). A comparison of the bonded geometry parameter distributions is shown in Figure SP9.
@@ 3.3 Nonbonded properties
The NB properties include the transverse strength and toughness, as shown in Figure 7. Table 4 lists the NB properties numerically. The reference strengths of the vertical, horizontal and slant models are 0.9 GPa[17], 0.5 GPa[17] and 0.7 GPa[38], which were well reproduced by both AA and CG simulations. To reduce the effect of randomness, 100 replica simulations of transverse stretch were performed. The strength and toughness distributions of the CG models were similar to those of the AA reference data. Only the distributions of the horizontal CG models were narrower than those of their AA counterparts were. The average strength and toughness values were similar to those obtained from the AA data. The frictional sliding of the slant models was quantitatively validated by using the direction angle (Figure SP9). These data validate the CG model for reproducing the transverse anisotropy and frictional sliding. Although all $\sigma$ values were larger than expected, the equilibrium distances in the CG models were only slightly larger than those in the AA models (less than %15, Table 2). A comparison of the nonbonded equilibrium distance distributions is shown in Figure SP9. The stretch curves (Figure SP9), radical distribution function, radius of gyration, and root-mean-square deviation were reproduced, as shown in Figure LABEL:fig:structural_properties.
@@
Vertical
Horizontal
Slant
Vertical
Horizontal
Slant
Figure 7: Nonbonded properties. (a) Behaviors of CG models in the characteristic directions. Important frictional sliding in the slant direction and fragile fractures in the vertical and horizontal directions are reproduced. (b) Strength and (c) toughness distribution of transverse stretches. A total of 100 replica simulations were performed with different random seeds for both AA and CG models in three directions. For the vertical and slant models, the strength and toughness histograms were reproduced consistently, because their mean values and standard errors were similar. Despite the accuracy of the mean values, the horizontal CG models provided narrower distributions, which was within expectations.
3.4 Additional inspection

### v3

@@ 2.3 Boltzmann inversion for harmonic potential
However, the $k$ derived must be reparameterized to represent the mechanical performance[74, 75, 49] of the CG models with fewer degrees of freedom. The effective $k$ is $k_{0}$ times of $\kappa$:
@@ 2.4 Reinforcement Learning and degenerate RL
In this study, a degenerate version of RL was leveraged in a one-shot diagram (Figure 3). There is only one step in a learning episode, and only one reward is sampled, because the performance of a set of coefficients is independent of that of the others. Therefore, the reward can be simplified as
@@
In this study, the RL agent was implemented using the Soft Actor Critic (SAC) algorithm[77]. SAC is the state-of-the-art algorithm recommended for continuous action space by the doc of OpenAI Stable-Baselines3[78]. The key feature is the reconciliation between the expected reward and entropy, which implies that stochasticity is included in the SAC policy to a certain extent. SAC is a gradient-free and model-free method that does not rely on a specific problem environment, and can address strongly nonlinear problems after sufficient training. Detailed information on the SAC algorithm can be obtained in the doc of OpenAI Stable-Baselines3[78].
During training, the action consists of 17 normalized coefficients ranging from [-1,1] because most RL algorithms rely on standard normal distributions[78]. The actions are linearly transformed to a physical space, which corresponds to 3 BD $k$, 1 HB $\epsilon$, 1 HB $\sigma$, 6 LJ $\epsilon$ and 6 LJ $\sigma$. The reward function is a critical factor that guides the learning of the agent, and is defined by a series of matching degrees. The matching degree $m$ for a specific performance is defined as
@@
. where $y_{0}$ and $y_{1}$ are the reference and sampled values, respectively. The reward function $R$ is the sum of a series of matching degrees, except for the transverse strength $M_{s}$ and toughness $M_{t}$,
$R=\sum M_{i}+w\cdot{\rm min}(M_{s},M_{t})$ (17)
@@ 2.5 Simulation setup
For both all atom (AA) and coarse-grained (CG) simulations, three models in different arrangement directions under the NPT ensemble at 0.1 MPa with a stretch load in the vertical direction were simulated. All reference AA simulations were performed using GROMACS[80, 81] and CHARMM36[82, 83] force fields (force field files were generated using CHARMM-GUI tools[84, 85]). All the CG simulations were performed using LAMMPS[69] with a self-defined force field. Referring to the performance-size curves obtained by Sinko et al.[86], we used a CNC model comprising 144 cellulose chains. Fragile vertical and horizontal models with 72 cellulose chains were used to reduce the computational cost of the simulations during training. Vertical stretches were implemented using cell deformation loads at a constant speed 10.0 nm/ns. 100 replica AA simulations with different stretch speed ranging from 0.1 nm/ns to 40.0 nm/ns were performed. The AA simulation results confirmed that the sampled strength and toughness were stable within this speed range (with deviations of less than 15%, as shown in Figure Supplementary Information). Figure Supplementary Information shows the stress-strain curves of the 100 replica simulations at 10.0 nm/ns. The timestep used in the CG simulations was 4 fs for the BD properties and 12 fs for the NB properties (transverse stretching), chosen as a result of bisect between 10 fs and 15 fs. The validation for the CG timestep was provided by a 12 ns CNC equilibrium thermodynamics under NPT ensemble using 12 fs timestep Supplementary Information.
@@ 2.6 Training procedures and statistics-guided reduction
Before training, the equilibrium geometry parameters (bond length $l^{e}$/,angle size $\theta^{e}$/, and dihedral size $\chi^{e}$) were derived from the mapped AA trajectories. The estimated force constants for the bonded harmonic interactions were determined using the same trajectories and Boltzmann inversion method. Three rescale factors were required for the bonds, angles, and improper dihedrals in this model. Although rescaled, the relative ratios of the force constants were maintained (e.g., the force constants for improper dihedrals were rescaled with the same factor). As previously mentioned, there are 17 coefficients, where three are BD coefficients and 14 are NB coefficients. To simplify the problem and reduce computational cost, three training stages were designed (Figure 4). The first stage was BD training for three BD $k$ only, ignoring all NB interactions and targeting the BD properties (axial elastic modulus and polymer stiffness). We cannot determine the influence of NB interactions in the first stage, and we assume that they are not critical for the BD properties. The second stage was NB training for 14 NB coefficients with the BD coefficients obtained from the first stage, targeting the BD and NB properties (transverse strength and toughness in the three characteristic directions). The third stage was preserved only if the performance in the second stage was unsatisfactory. All properties were targeted to parameterize 17 coefficients at every turn.
@@
Figure 4: Training procedures. The BD properties (axial elastic modulus and polymer stiffness) and NB properties (transverse strength and toughness in characteristic directions) were the target attributes. Before training, the equilibrium bonded geometry parameters and force constants of the BD harmonic potentials were estimated from the mapped AA trajectories using the Boltzmann inversion method. In the first stage, the force constants were scaled by the RL agent in the BD training to reproduce the BD properties. In BD training, all NB interactions were ignored because we assumed that the NB interactions were not significant in this case. Given both BD and NB properties, NB training using the previously obtained BD coefficients was performed to match both BD and NB properties. However, only one batch of simulations was performed for the target properties during training. Thus, the transverse stretch performance data were manually verified through replica simulations to confirm the statistical agreement. Multiple sets of coefficients were collected to reduce the number of independent coefficients for the computational cost and physical explainability. These results confirm that our assumption of the BD properties is reasonable. Therefore, the coupled training was unnecessary.
@@ 3.1 Coefficients
Figure 5: Training convergence and statistics-guided reduction. (a) Reward and its moving average in the first and (b) last training loops. Each training loop contained 1024 training steps. One trial took about 4 minutes and overall training took about 40 days. The subsequent training loops with higher and more stable rewards were based on the former training loops. (c) Average and maximum rewards for different training loops. The average and maximum rewards confirmed the convergence. After training, the best coefficients for the different training loops were collected to perform statistics-guided reduction. The first reduction was inferred from the geometric characteristics of the CG models, and a constraint of $\sigma_{p11}=\sigma_{p22}=\sigma_{p33}$ was applied (pmn represents a pair of CLm and CLn), which is also helpful for fragile behaviors in the vertical and horizontal directions. (d) The second reduction was in the distance coefficients using statistics. $\sigma$ except for $\sigma_{p23}$ oscillated around its equilibrium distance $d^{e}$ whereas an obvious increase of $\sigma_{p23}$ was observed. Then only $\sigma_{p11}$ and $\sigma_{p23}$ are kept as independent distance coefficients: $\sigma_{p12}=\frac{d^{e}_{p12}}{d^{e}_{p11}}\sigma_{p11},\sigma_{p13}=\frac{d^{e}_{p13}}{d^{e}_{p11}}\sigma_{p11}$. (e) The third reduction in energy coefficients was statistically determined. After the first and second reduction, we observe that $\epsilon_{p23}$ is usually the smallest energy coefficients, which may correspond to the $\sigma{p23}$ enlargement. Then only $\epsilon_{p23}$ and $\epsilon_{p23}$ are kept as independent energy coefficient: $\epsilon_{p11}=\epsilon_{p22}=\epsilon_{p33}=\epsilon_{p12}=\epsilon_{p13}$.
The coefficients were trained using several training loops, each containing 1024 training steps. When a training loop was completed, the information of the agent was saved and loaded for subsequent training. Considering the time consumption, 4 minutes for one trial and 3 days for a training loop was spent. The convergence criteria was defined as relative error between max reward of the latest episode and previous episodes. For the $n$-th training loop , the convergence criteria $C$ was defined as the relative advantages between max reward $R^{max}_{n}$ of this training loop and all of the former max reward $R^{max}_{1:n-1}$:
$C_{n}=\frac{R^{max}_{n}}{R^{max}_{1:n-1}}$ (19)
. If $C_{n}$ is lower than 0.05 for 5 training loops, the training is considered as convergent. Referring to the first training loop, the last training loop achieved a higher reward much faster and more stably (Figure 5a). The average and maximum rewards of the different training loops were compared (Figure 5b) to demonstrate the convergence of training. The component match degrees of reward function with learning steps were further provided in Figure Supplementary Information.
@@
where pmnpmn is the pair of the CLm and CLn. The structures of the CG models with lower degrees of freedom were more orderly. Thus, the equilibrium distances of the same types should be nearly identical in order to represent the laminar structure. In other words, the equilibrium distance of CL1-CL1, CL2-CL2 and CL3-CL3 ought to be nearly the same, whereas in the mapped AA trajectories, they are not (Figure Supplementary Information, Figure Supplementary Information). The equilibriumequilibrium distancesdistances of the ideal crystal structure were selected as the estimated value of the distance coefficient ($\sigma$) for each pair. These equilibriumequilibrium distancesdistances do not correspond to the well-relaxed structures or characteristic distances of the radial distribution function. The first physics-informed restraint was also beneficial for reproducing fragile fractures in vertical and horizontal models.
@@
These reductions and resultant coefficients informed us one thing: the sampled average values from AA trajectories could not be directly used as equilibrium points for CG interactions. The laminar structure and coefficients of CNC is critical to be artificial, and the $\theta_{0}$ was also chosen of the ”equilibrium angle” in the artificial model (the strict laminar model build by coordinate offset). This estimated $theta_{0}$ is beneficial but not strongly supported by molecular details.
@@
bond 1 64.06 5.270 5.270 5.258 5.262 5.263
2 82.42 2.493 2.493 2.473 2.493 2.479
3 112.3 2.066 2.066 2.061 2.066 2.065
kk(kcal/mol) $\theta_{0}(^{\circ})$ $\theta^{e}_{\rm AA,E}(^{\circ})$ $\theta^{e}_{\rm CG,E}(^{\circ})$ $\theta^{e}_{\rm AA,S}(^{\circ})$ $\theta^{e}_{\rm CG,S}(^{\circ})$
angle 1 30.97 163.5 163.5 162.3 163.7 161.7
2 38.79 74.2 74.2 74.7 74.2 75.4
3 30.52 93.2 93.2 94.0 93.3 93.4
4 27.28 90.0 90.0 89.2 90.2 89.9
5 25.97 102.0 102.0 101.4 102.0 100.1
6 22.42 166.3 166.3 165.1 166.5 165.1
kk(kcal/mol) $\chi_{0}(^{\circ})$ $\chi^{e}_{\rm AA,E}(^{\circ})$ $\chi^{e}_{\rm CG,E}(^{\circ})$ $\chi^{e}_{\rm AA,S}(^{\circ})$ $\chi^{e}_{\rm CG,S}(^{\circ})$
improper 1 0.1702 178.0 178.0 175.6 177.0 171.2
2 0.1555 2.3 2.3 2.1 3.2 3.5
3 0.1698 2.9 2.9 2.1 3.6 3.8
@@
type pair $\epsilon$(kcal/mol) $\sigma$(Å) $d^{e}_{\rm AA,E}$(Å) $d^{e}_{\rm CG,E}$(Å) $d^{e}_{\rm AA,S}$(Å) $d^{e}_{\rm CG,S}$(Å)
@@ 3.2 Bonded properties
For the BD interactions and validation of the BD coefficients, we considered the axial elastic modulus and polymer stiffness of a single cellulose chain, as listed in Table 3 and shown in Figure 6. All the values after $\pm$ are the corresponding standard error. Axial stretching was performed in an xyz-periodic cell with a deformation load along the chains, which is a basic mechanical property. The simulation of polymer stiffness is a single chain centered in a cell. For axial stretch, Poisson effect was presented in CG model and elastic modulus of AA and CG are 133.5 GPa and 130.3 GPa respectively, where reference modulus is 120-140 GPa[13, 14] from experiments and 110-200 GPa[16, 38, 17] from simulations. For polymer stiffness, persistence length of AA and CG range from 9.7-9.8 nm and 8.9-10.2 nm with reference data of 6-25 nm from experiments[87] and simulation[88]. Notably, the polymer stiffness data were not stable for the AA simulations because of the intertwining in some cases. Therefore, persistence lengths and end-to-end distances of different lengths (10, 20, and 30 residues) are listed. When considering polymer stiffness, the joint O atoms between two cellulose residues in the AA models and CL1 beads on the backbone in the CG models were node particles to be considered. The reference data for training were based on AA information only, and the mapped AA trajectories were compared with the CG trajectories after training. In this study, all effective $k$ values were smaller than their estimations based on distributions (Table Supplementary Information), particularly for angles and dihedrals, to reproduce the polymer stiffness performance in a system with fewer degrees of freedom (Table Supplementary Information).
@@ 3.3 Nonbonded properties
The NB properties include the transverse strength and toughness, as listed in Table 4. The transverse behavior of CG models are shown in Figure 7. The reference strengths of the vertical, horizontal and slant models are 0.9 GPa[17], 0.5 GPa[17] and 0.7 GPa[38], which were well reproduced by both AA and CG simulations. To reduce the effect of randomness, 100 replica simulations of transverse stretch were performed. The strength and toughness distributions of the CG models were similar to those of the AA reference data. Only the distributions of the horizontal CG models were narrower than those of their AA counterparts were. The average strength and toughness values were similar to those obtained from the AA data. The frictional sliding of the slant models was quantitatively validated by using the direction angle (Figure Supplementary Information). These data validate the CG model for reproducing the transverse anisotropy. Although all $\sigma$ values were larger than expected, the equilibrium distances in the CG models were only slightly larger than those in the AA models (less than 15%, Table 2). Although the accuracy of strength and toughness distributions is problematic owing to the representability problem of CG models, the full distributions of transverse strength and toughness are shown in Figure Supplementary Information for reference. Other stretch curves collections are illustrated in Figure Supplementary Information.
@@
Figure 7: Transverse behaviors of CG models in the characteristic directions. Important frictional sliding in the slant direction and fragile fractures in the vertical and horizontal directions are reproduced.
3.4 Structural properties
Structural properties including BD geometry parameter distributions, NB equilibrium distance distributions, and spatial correlations are felicitous quantitative measures to validate CG models. The bond and pair distribution between CL1 and CL1, and the radical distribution function (RDF) of the whole laminar CNC, within one HBonding layer, and inter HBonding layers are shown in Figure. The RDF of one layer and inter layer were computed from sliced equilibrium structures. The alignments for the spatial correlations confirm the effectiveness of this CG model. Full comparison of the BD geometry parameter distributions and NB equilibrium distance distributions in equilibrium state and stretching progress are illustrated in Figure Supplementary Information, Figure Supplementary Information, Figure Supplementary Information, and Figure Supplementary Information. Besides, radius of gyration (Rg), and root-mean-square deviation (RMSD) of AA and CG were shown in Figure Supplementary Information.
Equilibrium
Stretch
Equilibrium
Stretch
Whole
HBonding layer
Inter layers
Figure 8: Structural properties. (a) BD geometry parameter distributions between CL1 and CL1 in equilibrium state and stretching progress. (b) NB equilibrium distance distributions between CL1 and CL1 in equilibrium state and stretching progress. (c) RDF of a whole CNC and RDF of a HBonding layer slice or inter HBonding layers slice in equilibrium state. These data confirmed the accurate structural reproduction of CG model in both equilibrium and stretching states.
3.5 Baseline methods for coarse grained models
Except for the RL parameterized CG potential, baselines models were also tested in parallel. The simulations we used for AA and CG baselines were the slant model equilibrium without any external loads under NPT ensemble as we used for the RL parameterized model. The reference AA trajectory was simulated for 20 ns with timestep 2 fs.
Bottom-up baselines included IBI, RE, and FM and Top-down baseline MARTINI 3 were compared to illustrate the necessity of this hybrid-hybrid model. Bottom-up methods were utilized by the implementations by VOTCA[75, 89, 90]. RE is focused on NB interactions, and the BD interactions coefficients for the model were from IBI. IBI and FM models were fully independently trained 100 steps with timestep 1 fs. One trial of IBI or FM took about 8 minutes, and the 100 steps took 15 hours and 24 hours. Top-down baseline was MARTINI 3 cellulose potentials from Polyply reference[67, 91].
We felt sorry that FM to match mean forces did not produce a stable structure (it ”explored” in the energy minimization). The IBI and RE models were checked under NVT ensemble without velocity generalization because the velocity and pressure manipulation would also lead to unstable structures (the periodic chains rotated and translated). Thus polymer stiffness simulation results were provided but problematic because we were not sure about the effect of velocity generalization in these cases. The transverse stretch strength and toughness were also problematic without pressure control. MARTINI 3 cellulose potentials were more robust than the bottom-up baselines and simulated under NPT ensemble.
The comparison focused on axial elastic modulus and transverse mechanical properties, as listed in Table 5 and Table 6. The worse performances of bottom-up baselines were within expectations and aligned with their limitations described in the Introduction section. It should be emphasized that all the polymer stiffness data were all from simulations with velocity generalization under NVT ensemble, regardless of the NVT or NPT marks. Except for the numerical properties in the tables, structures after relaxation and fracture are illustrated in the Figure Supplementary Information and Figure Supplementary Information. Owing to its weak NB interactions and high fracture strain and toughness, fracture structures shown for MARTINI in the horizontal and slant characteristic directions was not fractured yet because it WAS stripped like a ribbon before fracture. The resultant results of these bottom-up baselines confirm the validation and superiority of this RL parameterized model following extended bottom-up approach.
Table 5: Baselines bonded properties
type Axial Modulus (GPa) Persistence Length (nm) End to end distance (nm)
10 20 30 10 20 30
IBI(NVT) 661.6 2.674 4.002 5.785 4.477 8.884 13.141
RE(NVT) 660.9 6.688 5.307 5.069 4.433 8.870 13.315
MARTINI(NPT) 10.7 5.440 5.054 6.807 4.134 8.078 11.907
RL(NPT) 130.5 8.990 10.176 10.106 4.529 9.199 13.432
AA(NPT) 133.5 9.755 9.815 9.716 4.523 9.355 13.536
Error
IBI(NVT) 396% 73% 59% 40% 1% 5% 3%
RE(NVT) 395% 31% 46% 48% 2% 5% 2%
MARTINI(NPT) 92% 44% 49% 30% 9% 14% 12%
RL(NPT) 2% 8% 4% 4% 0% 2% 1%
Table 6: Baselines nonbonded properties
type Strength (MPa) Toughness (GPa)
V H S V H S
IBI(NVT) 654.5 240.5 495.2 109.3 79.3 334.8
RE(NVT) 833.3 260.9 491.7 87.5 77.6 344.4
MARTINI(NPT) 471.6 268.1 259.7 62.0 396.6 838.7
[…]
Error
IBI(NVT) 47% 62% 6% 2% 2% 76%
RE(NVT) 32% 59% 5% 21% 0% 81%
MARTINI(NPT) 61% 58% 44% 44% 309% 341%
RL(NPT) 3% 2% 9% 5% 0% 12%
3.6 Training without thresholds and training using other optimization methods
The advantages of the thresholds and min in the reward function were proved a parallel training without thresholds and min function were provided, as shown in Figure Supplementary Information and Supplementary Information. The results from parallel training elucidated that the thresholds did not harm best performances and accelerate learning in the early stage.
On the other hand, the parameterization of interaction potentials is similar to the hyperparameter optimization of neural network models. The parameterization of interaction potentials is similar to the hyperparameter optimization of neural network models. To further validate the effectiveness of RL, parallel training using other optimization methods including Tree-structured Parzen Estimator[92] (TPE, an implementation of Bayesian Optimization) and Covariance Matrix Adaptation Evolution Strategy[93] (CMAES) were also performed, which are widely applied for hyperparameter tuning. CMAES and TPE was utilized by the implementation of OPTUNA[94], and both the reward function with/without thresholds are both trained (Figure Supplementary Information). The comparison emphasized the optimization capability of RL to deal with high nonlinear potential parameterization problems and keep critical stochasticity to compensate for insufficient sampling.
3.7 Additional inspections